# Identifying the Truth of Global Model: A Generic Solution to Defend Against Byzantine and Backdoor Attacks in Federated Learning (full version)

### v2

@@ Appendix 0.C More Results on Gaussian Noise Attacks
During the Gaussian noise attack, we chose not to combine it with the constrain-and-scale attack. This decision was based on the goal and implementation of the constrain-and-scale attack, which involves training two loss functions on the same model during each epoch with benign and adversarial datasets. Therefore, including the constrain-and-scale attack would not be helpful for the Gaussian noise attack, as the adversary would be directly modifying their model.
@@
Gaussian Noise Attacks on FMNIST: Figure 11 offers our results for the Gaussian-noise attack on the FMNIST dataset. We have grouped the aggregation algorithms into three categories based on their accuracy after the 100th iteration. The top performers among them were FedTruth, FedTruth-layer, and FLTrust with their final accuracy above $65\%$. The Median, Trimmed mean, and FedAvg algorithms comprise the second cluster, with a final accuracy of around $35\%$ to $45\%$. The final grouping of the lowest-performing algorithms under this attack configuration is FLAME and Krum, with a test accuracy of $30\%$ after the 100th iteration.
We first take a closer look at the lowest-performing group (FLAME and Krum). We presume the subpar performance for FLAME is most likely a result of the clipping effects, as this is the stage in the algorithm responsible for removing models whose weights seem to be scaled or boosted. However, with the number of adversaries (30% each round) and large non-iid sampling rate (80%), it seems to cause the clipping stage of the algorithm to lose its effectiveness over time. We suspect that Krum is having problems distinguishing the adversarial model cluster from the benign model clusters during this attack.
Interestingly, we observed that the Median, Trimmed mean, and FedAvg algorithms in the second group showed some peculiar behaviors, having initially performed similarly to the group with the highest-performing algorithms after 100 iterations; however, despite this finding, with a final accuracy between 35% to 45%, they demonstrated a high global model accuracy of 60 to 70% before gradually declining in performance from the 70th iteration. This decline persisted until the end of the experiment, indicating a significant decrease in performance for these two algorithms. It has come to our attention that the FedAvg algorithm is experiencing a negative impact from the noised models of the three adversaries. This effect has become apparent after several rounds of basic averaging without any assistance from boosting or scaling techniques. In contrast, the Median algorithm selects a model from a group of sorted models. As more adversarial models are added to the group, the probability of selecting one as the global model for a given iteration increases. Furthermore, due to the random distribution of the data, the likelihood of selecting an adversarial model also increases. The Trimmed mean algorithm combines the Median and FedAvg algorithms by removing outliers and averaging the remaining models, also producing a similar suboptimal performance.
@@
The group of high-performing aggregation algorithms after 100 iterations, which included FedTruth, FedTruth-layer, and FLTrust, produced the best results by using truth discovery to calculate weights for each adversarial model. However, we noticed that FLTrust experiences slight decreases in accuracy around the 30th, 50th, and 90th iterations. This may be due to the non-iid degree of the data sampling, causing the server model to not accurately represent the client’s model. As a result, the weight assignment may overweight the adversarial models too high. Thankfully, the adaptive truth discovery algorithm used by FedTruth and FedTruth-layer solves this problem and avoids any dips in accuracy.
Figure 11 also presents our results for the Gaussian Noise Attack with model-boosting attack when there are 3 adversaries per round and performing model boosting. Our results in Figure 11 can be divided into two groups based on their final accuracy: those algorithms that achieve an accuracy above 50% and those that fall below this threshold. The former group includes FedTruth, FedTruth-layer, FLTrust, and Median. The latter group, which consists of Trimmed Mean, FedAvg, and FLAME, Krum, experienced poor performance. Interestingly, the addition of model-boosting allowed the Median algorithm to achieve an overall accuracy of $70\%$, which is higher than during the base attack. This could indicate that the algorithm was better at distinguishing between adversarial and benign models after model-boosting.
Gaussian Noise Attacks on CIFAR-10: Figure 12 offers our results for the Gaussian Noise attack combined with and without the model-boosting attack on the CIFAR-10 dataset. During this attack, it is apparent that the FedTruth-layer is among the few algorithms that can reach convergence. Additionally, this attack configuration significantly hinders FedAvg, FLAME, and Trimmed mean. At the same time, Krum and Trimmed mean suffer a much slower convergence speed during this attack. The results for the base and model boosting attack combinations resulted in similar clustering of the aggregation algorithm. Excluding FedTruth and FedTruth-layer, all algorithms suffered a slight drop in accuracy after 1000 rounds when using the model-boosting attack combination. Furthermore, it is interesting to note that the Gaussian Noise attack with a model-boosting attack was more effective at degrading the performance of the CIFAR10 dataset than the FMNIST dataset. However, regardless of the attack combination or dataset selection, it is apparent that FedTruth and FedTruth-layer are consistently among the top performers.
@@ Appendix 0.D More Results on Backdoor Attacks
In this section, we present more finding for target task accuracy (accuracy on an adversarial backdoor dataset) and the main task accuracy (accuracy on a benign dataset) for the Projected Gradient Descent and Edge Case attacks observed during each round of aggregation.
Projected Gradient Descent Attack (PGD): We implemented the PGD attacks with the Torch Attacks [14] library, which creates a generative model that will take as input an image and return a perturbed version of the image. We set the max perturbation ($\epsilon=8/255$), which is how the adversarial example knows how far an image can be noised while generating the adversarial model. Then we set the step size ($\alpha=2/255$) and the number of steps ($4$).
Figures 13(a)14(a), and 15(a) present the main task accuracy for the PGD attacks. We observe that during the base (Figure 13(a)) attack, Krum and Flame do not reach convergence on the main task. During the model-boosting (Figure 14(a)) attack, only Krum is not able to reach convergence. With that being said, FedTruth, FedTruth-layer, and Flame have a slower convergence speed but still finish the 100th iteration with an accuracy above $70\%$. Figure 15(a) shows that during the PGD with constrain-and-scale attack, all aggregation algorithms are able to reach convergence with a slight decrease in the final accuracy of FedTruth after the 100th iteration to around $60\%$.
When evaluating the results on the targeted task accuracy, we did not consider any algorithms that could not converge on the main task (i.e., Krum and FLAME). Our results for the PGD attack (Figure 13(b)) and PGD with model-boosting attack (Figure 14(b)) show that FedTruth and FedTruth-layer were the only algorithms not excluded that exceed the $40\%$ backdoor accuracy threshold after the 70th iteration. However, when the PGD attack is combined with the constrain-and-scale attack, presented in Figure 15(b), seems to be even more effective as it is able to degrade the performance of FedAvg, Trimmed mean without effecting the main task accuracy.
@@
Figure 18: Edge Case Attack - combine with comtrain-and-scale attack (MNIST, 3 adversaries)
@@
Edge-case Attack: Based on the attacks presented in [29], we implemented the Edge Case attack for the MNIST dataset. During this attack, we used the Arkiv Digital Sweden (ARDIS) [17] dataset as adversarial images (edge-cases) being injected as backdoors into the models. During each training round, the adversaries examined their benign local data points to locate those corresponding to the targeted labels. Using this knowledge, each adversarial client added targeted edge-case data points to their training dataset. The number of these added data points was set to be equal to $20\%$ of the matching benign data points.
@@
Figures 16(a)17(a), and 18(a) show our results for the main task during the Edge-case attacks. We observed that all the algorithms could reach convergence during the base attack. However, during the model-boosting (Figure 17(a)) and constrain-and-scale (Figure 18(a)) attacks, FedTruth and FedTruth-layer had a slower convergence rate.
@@
Figures 16(b)17(b), and 18(b) present our results for the targeted task accuracy. During the base attack (Figure 16(b)), we see that the least effective algorithms were Trimmed mean and Median, with a final backdoor accuracy around $60\%$ after 100 iterations. FedTruth-layer, FedAvg, and FLTrust mostly removed the backdoor, with a final backdoor accuracy below $40\%$. FedTruth, Flame, and Krum were the most robust algorithms to this attack, finishing with a targeted task accuracy below $15\%$ after 100 iterations.
Using a threshold of $40\%$ targeted task accuracy after the 50th iteration to denote a successful attack combination, we see that FedAvg and FedTruth appeared to have difficulty removing the adversaries during the base attack, as seen in Figure 16(b). We observed FedTruth reaching a $60\%$ targeted task accuracy by the 100th iteration. However, we suspect these results occur due to the similarity between adversarial edge-cases and the benign dataset, along with the relatively small number of edge-cases
Figure 17(b) presents our results for the backdoor accuracy during the Edge-case with model-boosting attack. We observed during this attack that the FLTrust, Median, and FedAvg algorithms were the least effective at removing the backdoor with a backdoor accuracy above $60\%$ after 100 iterations. During this attack, Krum and Trimmed Mean performed much better, reducing the final backdoor accuracy to below $40\%$. However, FedTruth and FedTruth-layer were the most robust, removing the backdoors from the final model and finishing with a targeted task accuracy below $5\%$.
@@
The Edge-case with constrain-and-scale attack (Figure 18(b)) shows that FedTruth and FLTrust were the most robust algorithms, finishing with a final accuracy below $5\%$. FedTruth-Layer was slightly less effective during this attack, finishing with a targeted task accuracy around $25\%$. Trimmed Mean, FLTrust, Krum, and Median produced similar results to FedTruth-Layer, finishing with a backdoor accuracy below $40\%$. During this attack, FedAvg was the most susceptible algorithm, finishing with a backdoor accuracy of $95\%$
@@ Appendix 0.E More Results on Distance Functions
Figures 19, 20, and 21 compare the effects that different distance metrics (Euclidean distance, Manhattan distance, angular distance, cosine distance, and custom distance) have on the FedTruth and FedTruth-layer. We compare these results during the Gaussian Noise (Figure 19), PGD (Figure 20), and Edge Case attacks (Figure 21). Additionally, combined the Byzantine (Gaussian noise) attack with model boosting and the backdoor (PGD and Edge Case) attack with model-boosting and constrain-and-scale attacks.
@@
Gaussian Noise Attack - Comparison of Distance Functions: Figure 19 presents our results for the efficiency of different distance metrics during the Gaussian noise attack with and without model-boosting. During both attacks, the algorithms that computed the difference between two points (Euclidean and Manhattan distances) performed the best with a final accuracy above $90\%$. Additionally, the FedTruth-layer with custom distance (combination of the distance between the angular and Euclidean distance) metric also performed over $90\%$ during this experiment. However, this was not the case for the FedTruth (custom distance), with the final accuracy being below $60\%$ during the base attack (Figure 19(a)) and below $40\%$ during the model boosting attack (Figure 19(b)). The remaining algorithms that are based on the difference between the vectors’ angle (angular distance and cosine distance) are not able to reach convergence during both attacks, with them reaching a final accuracy below $20\%$.
This is consistent with our expectation, as adding arbitrary noise would not change the model enough to be represented by the difference between the two vectors when there are three adversaries present per round. This is exasperated when combining the Gaussian noise attack with model-boosting as it would cause the adversaries to degrade the performance of the models using the difference between the vectors’ angles. Furthermore, we suspect that the loss of precision in the FedTruth (custom distance) approach allows the inconsistent distance from the vectors’ angles computation to degrade its performance but not the FedTruth-layer (custom distance) due to its granularity.
@@
PGD - Comparison of Distance Functions: Figure 20 presents our findings for the PGD base, with model-boosting, and with constrain-and-scale attacks. Figures 20(a)20(c), and 20(e) illustrate how the PGD attacks impacted the main task accuracy during these experiments. We observed through these results that all of the algorithms are able to reach convergence during all attack combinations, except for FedTruth-Layer (angular distance) during the base attack. We additionally observe a slight decrease in the convergence speed of FedTruth and FedTruth-layer for the cosine distance distance metric when running this experiment on the model-boosting version of the attack, as seen in Figures 20(c). However, after the 100th iteration, the main task accuracy increases back to above 70% for the FedTruth (cosine distance) algorithm and above 80% during the FedTruth-Layer (cosine distance) versions. Figure 20(e) shows that during the PGD with constrain-and-scale attack, we observe that FedTruth (cosine distance) results slightly degrade after the 100th iteration to around $60\%$.
Figures 20(b)20(d), and 20(f) show our results for the targeted task accuracy during the PGD attacks. When considering the base attack (Figure 20(b)) and the PGD with model-boosting attacks (Figure 20(d)), all combinations of FedTruth and FedTruth-Layer successfully eliminated the backdoor artifacts across all distance metrics. After the first 20 iterations, none of the algorithms had a backdoor accuracy over $20\%$, and they all reached a backdoor accuracy below $5\%$ after 100 rounds.
Figure 20(f) presents our results during the PGD with constrain-and-scale attacks, where all of the attacks that use the difference between the angles (i.e., angular distance, cosine distance, custom distance) were less effective at removing the backdoor accuracy during training. We suspect that this is due to the model scaling changing the magnitude of the model weights significantly, while the angular distance of the adversarial models is small enough not to be excluded during training. Consequently, the metrics that do not depend on the angle of the model were able to account for adversarial model scaling properly. As a result, Euclidean and Manhattan distances effectively identified and suppressed backdoor models, reducing targeted task accuracy below $20\%$ by the end of the experiment. However, after 100 iterations, FedTruth and FedTruth-Layer using cosine distance successfully eliminated the backdoor artifact, achieving a final targeted accuracy of less than $5\%$. In contrast, FedTruth with the angular distance metric failed to remove the backdoor, as the targeted task accuracy reaches $100\%$ throughout the attack.
@@
Edge-case Attack - Comparison of Distance Functions: Figure 21 presents our findings on the impact of different distance metrics for both FedTruth and FedTruth-layer on the Edge Case attack.
Figures 21(a)21(c), and 21(e) present our findings for main task accuracy during the Edge Case attacks. We observe that all attacks reach convergence, with a final accuracy above $70\%$ across all attack configurations. However, the convergence rate is significantly slower during the Edge Case attack with model-boosting (Figure 21(c)) and constrain-and-scale (Figure 21(e)) attacks when FedTruth and FedTruth-Layer use cosine distance and angular distance metrics.
Since we are only considering a few edges (images from the ARDIS dataset), the adversarial models??? angle does not deviate significantly during the early stages of training. This makes the Edge Case with model-boosting and constrain-and-scale attacks more effective in degrading the performance of models using cosine distance and angular distance. Despite this, as training continues, they still reach convergence after the 100th iteration.
Figures 21(b)21(d), and 21(f) show our results for targeted task accuracy during the Edge Case attacks. Across all versions of the attack, the cosine distance metric consistently performs among the best for both FedTruth and FedTruth-Layer algorithms.
Figure 21(b) shows that during the base attack, FedTruth (cosine distance) achieves the lowest backdoor accuracy, falling below $15\%$, while FedTruth-Layer (cosine distance) maintains a final backdoor accuracy below $40\%$. Figures 21(d) and 21(f) present results for the Edge Case attacks with model-boosting and constrain-and-scale. These results indicate that cosine distance is the most effective distance metric, as the targeted task accuracy drops below $5\%$ after 100 iterations in the model-boosting attack and below $20\%$ in the constrain-and-scale attack.
Additionally, FedTruth (angular distance) produced results comparable to those of FedTruth (cosine distance) and FedTruth-Layer (cosine distance) in both attacks. However, this was not the case for FedTruth-Layer (angular distance), which performed among the worst, with backdoor accuracies exceeding $80\%$ after 100 iterations for both attack variations.

### v3

@@ Appendix 0.C More Results on Gaussian Noise Attacks
Figure 11 presents our results for the Gaussian noise attack with and without model-boosting. We did not combine this attack with the constrain-and-scale method, whose design trains both a benign and adversarial model each epoch using separate loss functions for each dataset. Therefore, incorporating constrain-and-scale would not be possible with the Gaussian noise attack, since adversaries aim to degrade model performance by directly modifying model weights and thus do not rely on an adversarial dataset.
@@
Gaussian Noise Attacks on FMNIST: Figure LABEL:fig:gaussian-attacks-fmnist-base offers our results for the Gaussian-noise attack on the FMNIST dataset when 3 adversaries are present. The top performers among them were FedTruth, FedTruth-Layer, and FLTrust with their final accuracy above $65\%$. The second-best clustering of algorithms, consisting of the Krum, FLAME, FedAvg, and Median algorithms, yielded final accuracies ranging from $45\%$ to $60\%$. The lowest performing algorithm during this experiment was Trimmed mean with a final accuracy of $40\%$.
We first examine the lowest-performing algorithm, Trimmed mean, which we hypothesize underperforms due to its tendency to select adversarial models during aggregation. This issue likely arises as a result of the added adversarial noise being constrained, which creates smaller differences between benign and adversarial updates, increasing the probability of the algorithm inadvertently selecting a malicious update during aggregation. Furthermore, compared to similar algorithms like FedAvg, Trimmed mean’s vulnerability to adversarial selection is exacerbated by its use of a subset of the models during aggregation. This limited subset of models heightens the attack’s effectiveness, as even minor adversarial influences disproportionately impact the final global model.
The cluster of models with the second-best performance (specifically Krum, FLAME, FedAvg, and Median) showed peculiar behavior, initially performing similarly to the highest-performing algorithms (reaching accuracies ranging from 60% to 70%); however, their final accuracy declined to between 45% and 60%. This decline persisted throughout the experiment, indicating a significant decrease in performance for these algorithms. We hypothesize two main reasons for this behavior. First, the constrained amount of noise often fails to differ substantially from the updates of a benign model during certain rounds, causing adversarial updates to be selected inadvertently. This occurs because the differences between adversarial and benign models are initially more pronounced due to the non-iid nature of the data. However, as the global model starts to represent the overall data distribution better, these differences become subtler, gradually leading to performance degradation. Second, the limited adversarial noise introduced into the models is diluted through the averaging process over all updates (as in FLAME and FedAvg), reducing the adversarial update’s overall impact; however, as this continues, it gradually reduces the performance of the global model during the later iterations.
@@
The high-performing aggregation algorithms (FedTruth, FedTruth-Layer, and FLTrust), used truth discovery to calculate weights for each adversarial model. However, we noticed that FLTrust experiences slight decreases in accuracy around the 70th iterations. This may be due to the non-iid degree of the data sampling, causing the server model to not accurately represent the client’s model. As a result, the weight assignment may overweight the adversarial models. Thankfully, the adaptive approach used by FedTruth and FedTruth-Layer solves this problem and avoids any dips in accuracy.
Figure LABEL:fig:gaussian-attacks-fmnist-model-boosting presents our results for the Gaussian Noise Attack when combined with the model-boosting attack. However, during this experiment, we observed results similar to those of the base attack (Figure LABEL:fig:gaussian-attacks-fmnist-base), with the top-performing cluster still being FedTruth, FedTruth-Layer, and FLTrust. Between the 30th and 60th iterations, FedTruth experiences a decrease in accuracy, reaching 40%. This dip does not persist in the final accuracy. The second-best cluster of algorithms, achieving final accuracies ranging from 50% to 60%, now consists of Median, Krum, and Trimmed Mean, all of which perform better during this attack than during the base attack. This improvement was anticipated, as boosting the noised updates makes the adversarial updates more apparent, allowing these aggregation methods to more reliably remove adversarial updates before selecting a representative model (Median and Krum) or averaging a subset of models (Trimmed Mean). FLAME and FedAvg become the worst-performing algorithms, finishing with accuracies of 38% and 30%, respectively. This outcome is expected, as the model-boosting attack can counteract the effects of adversarial weight delusion present during the base attack, making the attack more effective.
Gaussian Noise Attacks on CIFAR-10: Figure 12 offers our results for the Gaussian Noise attack combined with the model-boosting attack (Figure LABEL:fig:gaussian-attacks-cifar10-model-boosting) and without (Figure LABEL:fig:gaussian-attacks-cifar10-base) on the CIFAR-10 dataset. During these attacks, it is apparent that FedTruth and FedTruth-Layer are among the best-performing algorithms. Additionally, these attack configurations significantly hinder FedAvg, Krum, and Trimmed mean. At the same time, FLAME and FLTrust experience a slower convergence rate. Additionally, during the base attack, FLAME has a backdoor accuracy of below 15%, while FLTrust exhibits unstable accuracy, reaching a value below 10% at the 40th and 80th iterations during the Gaussian noise attack with model-boosting. Excluding FedTruth and FedTruth-Layer, all algorithms except for the Median algorithm suffered a slight drop in accuracy after 100 rounds when using the base or model-boosting attack combination. Furthermore, it is noteworthy that the Gaussian Noise attack combined with a model-boosting attack was more effective at degrading the performance of the CIFAR10 dataset than the FMNIST dataset. However, regardless of the attack combination or dataset selection, it is apparent that FedTruth and FedTruth-Layer are consistently among the top performers.
@@ Appendix 0.D More Results on Backdoor Attacks
In this section, we present more findings for target task accuracy (accuracy on an adversarial backdoor dataset) and the main task accuracy (accuracy on a benign dataset) for the projected gradient descent and edge-case attacks observed during each round of aggregation. We used the cosine distance metric during these experiments when running the FedTruth and FedTruth-Layer algorithms.
[…]
Figures 13(a)14(a), and 15(a) present the main task accuracy for the PGD attacks. During all of the attack, Krum and Flame do not reach convergence on the main task. During the constrain-and-scale (Figure 15(a)) attack, FedTruth and FedTruth-Layer have a slower convergence rate than the other models that were able to reach convergence. However, these algorithms still reach a final accuracy above 80%.
Figures 13(b)14(b), and 15(b) show our results for the targeted task accuracy during the attack combinations considered in this section. Krum and FLAME failed to converge on the main task and were excluded from the targeted task accuracy analysis. The results indicate that of the algorithms that demonstrated convergence on the main task, FedAvg was the only algorithm to exhibit significant vulnerabilities to backdoor attacks, achieving a final backdoor accuracy above 80% across all tested attack scenarios.
@@
Figure 18: Edge Case Attack - combine with constrain-and-scale attack (MNIST, 3 adversaries)
@@
Edge-case Attack: Based on the attacks presented in [29], we implemented the edge-case attack for the MNIST dataset. During this attack, we used the Arkiv Digital Sweden (ARDIS) [17] dataset as adversarial images (edge-cases) being injected as backdoors into the models. During each training round, the adversaries examined their benign local data points to locate those corresponding to the targeted labels. Using this knowledge, each adversarial client added targeted edge-case data points to their training dataset. The number of these added data points was set to be equal to $20\%$ of the matching benign data points.
@@
Figures 16(a)17(a), and 18(a) show our results for the main task during the edge-case attacks. We observed that all the algorithms could reach convergence during the base attack. However, during the model-boosting (Figure 17(a)) and constrain-and-scale (Figure 18(a)) attacks, FedTruth and FedTruth-Layer had a slower convergence rate than the other algorithms, excluding FedAvg, which had the slowest convergence rate during all of the attack configurations.
@@
Figures 16(b)17(b), and 18(b) present our results for the targeted task accuracy. During the base attack (Figure 16(b)), we see that the least effective algorithm was Trimmed mean with a final backdoor accuracy around $40\%$ after 100 iterations. The remaining algorithms, FedTruth-Layer, FedAvg, Krum, and FLTrust, mostly removed the backdoor, with a final backdoor accuracy below $20\%$.
Figure 17(b) presents our results for backdoor accuracy during the edge-case with model-boosting attack. We observed that FLTrust and Krum were the best-performing algorithms, both finishing with a backdoor accuracy below 20%. The Median and FedAvg algorithms also performed well, finishing with a backdoor accuracy of slightly above 25%. However, Trimmed Mean, FedTruth, and FedTruth-Layer did not perform as well, each finishing with a backdoor accuracy around 60%. We suspect that the degraded performance of our algorithm is due to the small difference between the adversarial and benign datasets, which allows adversarial updates to receive too high a weight during aggregation. As a result, when model boosting is applied, the adversarial model can hijack the global model and insert the backdoor artifact. Nevertheless, when using a different distance metric that takes into consideration the magnitude of the client updates (i.e., Euclidean Distance), we observe better performance, as discussed in Appendix 0.E.
@@
Figure 18(b) shows the effect of the edge-case attack with constrain-and-scale where FedTruth, FLAME, and Krum are the most robust algorithms, finishing with a final accuracy below 10%. FedTruth-Layer was slightly less effective during this attack, finishing with a targeted task accuracy of around $30\%$. FLTrust, Krum, and Median produced similar results to FedTruth-Layer, finishing with a backdoor accuracy below $40\%$. During this attack, Trimmed Mean and FedAvg were the most susceptible algorithms, finishing with a backdoor accuracy of $60\%$.
@@ Appendix 0.E More Results on Distance Functions
Figures 19, 20, and 21 compare the effects that different distance metrics (Euclidean, Manhattan, angular, cosine, and custom distance) have on the FedTruth and FedTruth-Layer. We compare these results during the Gaussian Noise (Figure 19), PGD (Figure 20), and edge-case attacks (Figure 21). Additionally, combined the Byzantine (Gaussian noise) attack with model boosting and the backdoor (PGD and Edge-case) attack with model-boosting and constrain-and-scale attacks.
@@
Gaussian Noise Attack - Comparison of Distance Functions: Figure 19 presents our results for the effectiveness of different distance metrics during the Gaussian noise attack with and without model-boosting. During both attacks, metrics based on vector magnitude differences (Euclidean and Manhattan distances) consistently outperformed the other approaches, maintaining a final accuracy above $90\%$.
In contrast, metrics based on the vectors’ direction, cosine, angular, and custom distances performed poorly during both experiments. Throughout the base attack, cosine, angular, and custom distance metrics maintained accuracies below $5\%$. During the model-boosting attack, the custom distance performed slightly better, with accuracy converging to around $40\%$, while angular distance converged to about $30\%$, and cosine distance remained low, converging to only around $5\%$.
This is consistent with our expectations, as adding arbitrary noise changes the magnitude of the adversarial updates, which is reflected in the strong performance of the Euclidean and Manhattan distances. However, these attacks do not significantly alter the angular difference when there are three adversaries present per round. We suspect that when the Gaussian noise attack is combined with model-boosting, the boosted updates make the custom distance function more effective at distinguishing between adversarial and benign updates, causing it to weight the adversarial updates lower. Nevertheless, the directional components in our custom metric ultimately limit its performance, causing the model to converge to a suboptimal global model.
@@
PGD - Comparison of Distance Functions: Figure 20 presents our findings for the PGD base, with model-boosting, and with constrain-and-scale attacks. Figures 20(a)20(c), and 20(e) illustrate how the PGD attacks impacted the main task accuracy during these experiments. We observed through these results that all of the algorithms are able to reach a final accuracy above 80% during all attack combinations, except for FedTruth (custom and angular distance) during the base attack. We observed a slight decrease in the convergence speed of FedTruth for the Euclidean and cosine distance metrics when running this experiment on the constrain-and-scale version of the attack, as seen in Figures 20(e). However, after the 100th iteration, the main task accuracy increases back to above 80% for these metrics.
Figures 20(b)20(d), and 20(f) present our findings on the effect of different distance metrics during the PGD attacks. Across all attacks, the cosine distance metric successfully eliminated the backdoor artifacts, reducing the final backdoor accuracy to under $5\%$.
During the base attack (Figure 20(b)), Euclidean and custom distances both performed poorly, with backdoor accuracy remaining close to $100\%$ throughout training. Manhattan distance also struggled, reaching a final backdoor accuracy of $80\%$. In contrast, angular distance was more effective, ultimately reducing backdoor accuracy below $20\%$. For the model boosting attack (Figure 20(d)), Euclidean distance again failed to suppress the backdoor, with a final accuracy close to $100\%$. Custom distance showed improvement in this setting, lowering backdoor accuracy to below $20\%$. Manhattan distance achieved moderate results, ending at $40\%$ backdoor accuracy. Angular distance also performed well, though it required more iterations to converge, ultimately reaching a backdoor accuracy below $5\%$. In the constrain-and-scale attack (Figure 20(f)), both Euclidean and Manhattan distances proved effective, each reducing backdoor accuracy to below $5\%$. The custom and angular distance metrics were ineffective, with a final backdoor accuracy near $100\%$.
Overall, these results demonstrate that cosine distance is consistently robust during all of these attack combinations, reliably reducing backdoor accuracy to below $5\%$. Euclidean and Manhattan distances are only effective when PGD is combined with the constrain-and-scale attacks. This highlights the importance of selecting an appropriate distance metric based on the specific attack scenario. We suspect that the strong performance of Euclidean and Manhattan distances during the PGD with constrain-and-scale attack is due to the model constraining-which slightly reduces the effectiveness of the model-combined with scaling, which makes the magnitude of the model easier to distinguish, thereby allowing these metrics to perform optimally. This effect is somewhat reflected in the model boosting results as well, where Manhattan distance is able to partially remove the backdoor and custom distance achieves an optimal global model.
@@
Edge-case Attack - Comparison of Distance Functions: Figure 21 presents our findings on the impact of different distance metrics for both FedTruth and FedTruth-Layer during the edge-case attack.
Figures 21(a)21(c), and 21(e) present our findings for main task accuracy during the edge-case attacks. We observe that all attacks reach convergence, with a final accuracy above $80\%$ across all attack configurations. However, the convergence rate is significantly slower in all of these attacks when using the cosine distance metric.
Figures 21(b)21(d), and 21(f) show our results for targeted task accuracy during the edge-case attacks. Figure 21(b) shows that during the base attack, the cosine distance and Euclidean distance metrics achieve the lowest backdoor accuracy, both falling below $15\%$, while FedTruth-Layer (cosine distance) maintains a final backdoor accuracy below $30\%$. Figure 21(d) presents our results for the edge-case attacks with model-boosting, where we see the final backdoor accuracy for cosine distance increase to 60%. This is consistent with our findings for the DBA attack when combined with model-boosting, as seen in Section LABEL:app:more-distance-functions-dba, where we suspect that the minimal change in the adversarial vectors’ angles allows boosting to be effective at hijacking the global model. Euclidean distance still offers the best performance during this attack, finishing with a final backdoor accuracy of 20%. During the constrain-and-scale attack (Figure 21(f)), the best performing metric is once again cosine distance, finishing with a backdoor accuracy of 30%, well below the other approaches. Additionally, Euclidean distance finishes with a backdoor accuracy of approximately 60%. We suspect this is a result of constraining the amount the model can change during aggregation, causing the model to alter the angular difference enough to weight the adversarial models low enough to remove them during training.